OpenStudy (anonymous):

What is the equation of a parabola with a directrix x = 8 and focus (-8, 0).

OpenStudy (agent47):

give me a second, reading up on how to do this.

OpenStudy (anonymous):

k

OpenStudy (agent47):

Let's follow along this example.

OpenStudy (agent47):

Let (x,y) be any point on the parabola. Find the distance between (x,y) and the focus, (-8,0): sqrt((x+8)^2+(y-0)^2) = sqrt((x+8)^2+y^2). The distance between (x,y) and the directrix, x = 8, is: |x-8|. Equate the two distance expressions and square both sides: sqrt((x+8)^2+y^2) = |x-8| (x+8)^2 + y^2 = (x-8)^2

OpenStudy (agent47):

wait let me double check everything.

OpenStudy (anonymous):

cool

OpenStudy (agent47):

ok, now I'm lost lol.. @myininaya @TuringTest @dpaInc does any of you remember how to do this?

OpenStudy (agent47):

*do

OpenStudy (agent47):

@Directrix

OpenStudy (agent47):

@jim_thompson5910

OpenStudy (agent47):

Can you bump this?

jimthompson5910 (jim_thompson5910):

The vertex is the midpoint of the segment that is perpendicular to the directrix and that goes through the focus So the vertex has a y coordinate of 0 The x coordinate of the vertex is (8 + (-8))/2 = 0/2 = 0 which means that the vertex is at (0,0)

OpenStudy (anonymous):

not yet.. in 15 min :/

jimthompson5910 (jim_thompson5910):

Now use the form 4p(x-h) = (y-k)^2 where p is the distance from the focus to the vertex (along the line of symmetry) and (h,k) is the vertex

OpenStudy (anonymous):

since the directrix is a vertical line, this means the parabola will have the equation: \(\large (y-k)^2=4a(x-h) \) where (h,k) is your vertex....